OpenStudy (anonymous):

The sum of the page numbers on facing pages of a book is 157; what are the page numbers?

OpenStudy (anonymous):

it wants to know an equation

OpenStudy (anonymous):

I think I am going wrong.. What are the facing pages do you know??

OpenStudy (anonymous):

Translate the following problem into algebraic equations that what the teacher wants

OpenStudy (anonymous):

Sorry, I have no idea..

OpenStudy (anonymous):

A woman buys three pairs of jeans for $99.84, including tax. If each pair of jeans cost $32, what was the sales tax rate?

OpenStudy (anonymous):

Can you translate this into an algebraic equation

OpenStudy (anonymous):

I put 32x=99.84 but not sure if its right

OpenStudy (anonymous):

\[x + 3 \times 32 = 99.84\] Where, x is the sales tax rate. \[x + 96 = 99.84\]

OpenStudy (anonymous):

Subtract 96 both the sides and find x..

OpenStudy (anonymous):

great thanks